OpenStudy (anonymous):

At George Washington High School the ratio of graduating seniors that go on to college, as compared to those who do not, is 7 : 4. Last year 28 students went on to college. How large was the graduating class? @ashley_f97 @fabsam14

OpenStudy (anonymous):

7(4)/4(4) = 28/16 28+16= 44

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\psi \]